A new weight-reducing technique, consisting of a liquid protein diet, is currently under- going tests by...

  1. A new weight-reducing technique, consisting of a liquid protein diet, is currently under- going tests by the Food and Drug Administration (FDA) before its introduction into themarket. A typical test performed by the FDA is the following: The weights of a random sample of 8 people are recorded before they are introduced to the liquid protein diet. The 8 individuals are then instructed to follow the liquid protein diet for 3 weeks. At the end of this period, their weights (in pounds) are again recorded. The results are listed in the table.

8 persons are selected 1 2 3 4 5 6 7 8 Before taking the diet: 134 204 167 276 187 168 175 176 After taking the diet: 130 206 155 258 175 166 170 159

(a) Test whether the diet is effective, use ?=0.05.
(b) What is the P-value?
Answer #1
Mud = population mean diffrence of the matched pairs data Ho: Mud = 0 Ha: Mud > 0 a = .10 test statistic = to to = d/(sd/square root of n) to = 5/(1.58/square root of 5) to = 7.076 p = 1 - tcdf(-1E99, 7.076, df = 4) p = 0.001 Since p < a, you would reject the null hypothesis and be able to conclude that there is enough evidence that the mean weight after 3 weeks on the diet is less than the mean beginning weight.